A piston of 46 mm diam moves concentrically in a cylinder of 50 mm diam. The cylinder is filled with water and when the piston moves into the cylinder the water flows through the gap surrounding the piston. If the velocity of the piston is 75 mm/s relative to the cylinder, what is the velocity of the flow through the gap.

1
Expert's answer
2021-11-05T03:39:36-0400

"Q= A\times V"

Q=  75 mm/s

A= Area

V =Velocity of the flow

"V =\frac{Q}{A}=\frac{75}{\frac{\pi}{4}46^2}=" 45.12 "\frac{m}{sec}"